OpenStudy (anonymous):

Corinne has a job selling magazines. She earns $7.50 per hour plus 20% of the total amount of her sales. She also gets an allowance of $40 per week for gas. She knows her weekly earnings can be shown using the following expression:7.50h + 0.20s + 40 Part A: Identify a coefficient, a variable, and a constant in this expression. (3 points) Part B: If Corinne works for 25 hours and sells $300 in magazines, how much does she earn for the week? Show your work to receive full credit.(4points) Part C: If Corinne gets a raise and begins earning $9 per hour, would the coefficient, variable, or

OpenStudy (anonymous):

I already have part a.

OpenStudy (anonymous):

Part b) (25*7.50)+(0.2*300)+40=187.5+60+40=187.5+100=287.5

OpenStudy (anonymous):

thank you! so for part c it the variables, coefficients, and constants would be the same if she got a $9 dollar raise right?

OpenStudy (anonymous):

The variables would stay the same, I think the constants would too, but the first coefficient wouldn't since instead of 7.50, it would be 9.

OpenStudy (anonymous):

the first coefficient will be a constant.

OpenStudy (anonymous):

Really? Oh, yeah right. sorry. Then I think you are correct.

OpenStudy (anonymous):

well thank you very much! fan and medal!

OpenStudy (anonymous):

Thanks! First medal, and first fan.